### v1

# thanks: equal contributionthanks: equal contribution

Abstract
Nature uses elongated shapes and filaments to build stable structures, generate motion, and allow complex geometric interactions. In this Review, we examine the role of biological filaments across different length scales. From the molecular scale, where cytoskeletal filaments provides a robust but dynamic cellular scaffolding, over the scale of cellular appendages like cilia and flagella, to the scale of filamentous microorganisms like cyanobacteria which are among the most successful genera on Earth, and even to the scale of elongated animals like worms and snakes, whose motility modes inspire robotic analogues. We highlight the general mechanisms that couple form and function. Physical principles, such as classical elasticity and the non-reciprocity of active matter can be used to trace unifying themes linking these systems spanning about six orders of magnitude in length.
I Introduction
The interconnectedness between form and function in biology has been quantified at least since D’Arcy Wentworth Thompson’s influential work Thompson (1917). Because geometrical features influence and direct biological activity through physical forces, it seems worthwhile to explore the link between geometry and biological function. In this Review, we examine a simple geometrical shape –the filament – which is pervasive in the architecture of living organisms, and discuss the function of filamentous shapes and the physical advantages they offer.
Filamentous shapes permeate every length scale in the biological world. Natural evolution has found that the geometrical properties of long aspect-ratio shapes can be used in a multiplicity of useful ways. The biopolymers forming the cytoskeleton provide structural integrity to the cell and form an intracellular transportation system. Cytoskeletal networks generate cytokinesis Eggert et al. (2006), migration Paluch et al. (2006); Lämmermann et al. (2008), and apical contraction Martin et al. (2009), and even generate cytoplasmic streaming in the Drosophila oocyte Dutta et al. (2024).
At length scales of a few microns, cilia and flagella can generate fluid flow and active motion due to their fluid dynamical interactions Lighthill (1976); Gilpin et al. (2020); Lauga and Powers (2009) and they can synchronize to generate metachronal waves Vilfan and Jülicher (2006); Gilpin et al. (2020). Microorganisms such as filamentous fungi form complex networks of hyphae following available water and resources, and can move cytoplasm along the network Klein and Paschke (2004). Filamentous fungi are also the group of fungi most often used in biotechnological processes and include perhaps the most important fungus in the history of humanity: Penicillium rubens, which was used by Alexander Fleming to extract penicillin Fleming (1929, 1941); Houbraken et al. (2011). The slime mold Physarum polycephalum performs cytoplasmic flow across its network akin to a generalised form of peristalsis, which is optimised when the wavelength of the peristaltic wave is of the order of the size of the network Alim et al. (2013); this filamentous network can even encode memory about nutrient location in its morphology Kramar and Alim (2021). In the prokaryotic domain, cyanobacteria were among the first microorganisms to straddle the transition from unicellular to multicellular organisms Grosberg and Strathmann (2007); Claessen et al. (2014), and include numerous filamentous species.
Box I | Geometry of filaments A filament can be described as a portion of a three-dimensional curve, whose local orientation changes smoothly along its length. The curve needs to be sufficiently smooth, that is, it is at least thrice continuously differentiable. The Frenet frame or trihedron is a local, right-handed, orthogonal coordinate system with basis $(\bm{t},\bm{n},\bm{b})$. We indicate the curve by $\bm{r}(s)$, where $\bm{r}$ is the position vector and $s$ is the arc length of the curve. The tangent unit vector to the curve is defined as $\bm{t}\equiv\frac{d\bm{r}}{ds}$. The normal unit vector $\bm{n}\equiv\frac{d\bm{t}}{ds}/|\frac{d\bm{t}}{ds}|$, and always satisfies $\bm{n}\cdot\bm{t}=0$. Furthermore, $\kappa\equiv|\frac{d\bm{t}}{ds}|$ is the local curvature of the curve $\bm{r}(s)$. Finally, the binormal vector $\bm{b}\equiv\bm{t}\times\bm{n}$ together with $\bm{t}$ and $\bm{n}$ form an orthonormal basis for space. These three vectors obey the Frenet–Serret equations $\frac{d\bm{t}}{ds}=\kappa\bm{n}\,,\quad\frac{d\bm{n}}{ds}=-\kappa\bm{t}+\tau\bm{b}\,,\quad\frac{d\bm{b}}{ds}=-\tau\bm{n}\,,$ where $\tau$ is the torsion. The Frenet trihedron can be useful in the description of the helical motion of free-swimming organisms Crenshaw and Edelstein-Keshet (1993); Friedrich and Jülicher (2009); Bente et al. (2020), sperm cells Jikeli et al. (2015); Friedrich and Jülicher (2007), chiral active particles Sevilla (2016), and for gliding motility Lettermann et al. (2024). [Uncaptioned image]
Figure 1: Filamentous shapes in biology. (a) Growing complexity of filaments across orders of magnitude in length, from nucleic acids that store information, to biopolymers that provide structural integrity to cells and form networks spanning the entire cytoskeleton. Flagella and cilia self-assemble from polymers to generate motion. Bacterial cells such as cyanobacteria self-organize into filamentous shapes. Animals, from worms and nematodes to snakes and eels utilize elongated shapes to achieve locomotion and collective behaviour. (b) Hypothetical schematic diagram of the biological function emerging from the three main properties of biological filaments: their length, their rigidity (measured by, e.g., Young’s modulus), and their structural complexity, such as twisting, cross-linking into a network, up until the tissue level of the muscle fibers or an entire body.
At much larger length scales, organisms such as worms, snakes, and eels have elongated bodies that grant them ecological advantages. Worms, for example, can form entangled ‘blobs’ that exhibit shear-thinning rheology Deblais et al. (2020) and reversible self-assembly Patil et al. (2023). Such advantages can be combined to create higher levels of complexity, such as in a network built from cross-linking filaments, as illustrated in Fig. 1.
Although the systems described above span about nine orders of magnitude in size (from nucleic acids in the nanometre range in diameter, to snakes and eels in the metre range), it is possible to identify and discuss general physical features of elongated shapes (from macromolecules to whole organisms) that depend on the geometrical form of filaments and, in turn, influence their biological activity.
How do cytoskeletal filaments generate structural integrity in the cell, how do groups of cilia generate fluid flow, or how do large elongated organisms generate undulatory locomotion? Understanding how physical systems interact via simple rules to generate interesting collective behaviour can and will have novel applications. The condensed-matter motto of ‘More is Different’ proposed by Philip Anderson Anderson (1972) is a guiding principle in this field. For example, a self-organizing robot exhibits behaviour intermediate between solid and liquid with adaptive capabilities such that it can flow over obstacles Saintyves et al. (2024). This combination of flexibility and structural integrity is reminiscent of a cell’s cytoskeleton, or even of snake locomotion. Each of these filamentous systems can serve as inspiration for new biosynthetic materials Fladung et al. (2024). Examples of this principle are the realisation of strain-stiffening metamaterials using PDMS slats that mimic cell behaviour Taale et al. (2023), or the use of carbon nanotubes to study the effect of crowding and filament stiffness on their microscopic dynamics Fakhri et al. (2010, 2009).
Physically, the rigidity and length of different biofilaments determine the class of models used to describe their mechanical and dynamical properties. If fluctuations are not important (because of the filament’s rigidity or because of the length scale of observation), the biofilaments can be effectively described as smooth lines. In this context, we can mathematically describe filaments as a curve in a three-dimensional (3D) space whose local orientation changes smoothly along their length. Box I describes the Frenet–Serret equations used in differential geometry to characterize a curve.
At smaller length scales, such as inside the cell, the polymeric nature of biofilaments is consequential and this character is amplified by the nonequilibrium nature of the fluctuations. For example, variations in polymer stiffness and connectivity of actomyosin networks by actin filament bundling and cross-linking proteins can systematically tune the mechanical response of actin networks from extensile to contractile Stam et al. (2017); Banerjee et al. (2020).
Another dimension, useful in classifying filaments in the natural world, is the complexity of their organisation. Single globular actin proteins (G-actins) assemble in filaments (F-actin), which together with cross-linking proteins and a myriad of other molecules form cytoskeletal networks. Multiple cells, in turn, can assemble into simple filaments, such as the Oscillatoria lutea cyanobacteria, or form even more complex filamentous organisms, such as nematodes, eels and snakes.
In this Review, we discuss the physical processes and properties of filamentous shapes in the biological world across length scales. Section II discusses the role of biofilaments inside the cell. Section  III reviews the role of filamentous cellular appendages, such as the locomotor function of cilia and flagella. In Sec. IV we discuss the biological properties of filamentous bacteria. Section V deals with macroscopic elongated organisms, such as snakes and eels. Finally, in Sec. VI we outline a roadmap of future work and outstanding questions at this exciting frontier shared by biology and physics, to establish a common framework describing the physics of biological filaments (Fig. 1)
II Biopolymers & cytoskeletal filaments
The cytoskeleton is a 3D polymer-based scaffold that spans the entire interior of a cell and defines cell morphology Huber et al. (2013). The basic building blocks are few key proteins that assemble together into a broad range of filaments Fletcher and Mullins (2010). These filaments together with a host of proteins that bind to the filaments’ sides or ends form a network that provides mechanical stability to the cell Pegoraro et al. (2017). Using only a few building blocks allows great flexibility and economy of materials. For example, cytoskeletal polymers achieve the same rigidity (elastic modulus) of gel-like substances with two orders of magnitude less material Kollmannsberger and Fabry (2011); Pegoraro et al. (2017).
Biological filaments are subject to different types of deformations: extensional and compressional forces and moments that cause the filaments to bend and twist. Any deformation mode costs energy, which is continuously used in the cytoskeleton to maintain the cell in a nonequilibrium state and allows dynamic control and reconfiguration Banerjee et al. (2020). The theory of elasticity can help identify the associated bending energy $E_{b}$ Landau et al. (2012), see Box II.
Generally, one expects the bending energy to be proportional to the square of the deformation, i.e., of the curvature; if the symmetry of the system does not distinguish positive and negative curvature, linear terms in the curvature cannot be present. To lowest order, the energy of deformation is then quadratic
[…]
which is known as the Kratky–Porod model Kratky and Porod (1949).
At the molecular level, cytoskeletal filaments are linear polymers: monomers that are bound together as an unbranched chain. Physically, they can be distinguished by their stiffness. Thermal fluctuations induce random deformations of the filaments; thus, the thermal energy $k_{\mathrm{B}}T$ is a natural scale to classify the ease of bending. The ratio $\lambda_{\mathrm{p}}=\frac{\beta_{\mathrm{f}}}{k_{\mathrm{B}}T}$ is the persistence length of the filament. By comparing $\lambda_{\mathrm{p}}$ with the filament length we can classify filaments as follows.
[…]
If $\lambda_{\mathrm{p}}\ll L$ the filament is considered flexible (the filament will exhibit highly coiled configurations); if $\lambda_{\mathrm{p}}\approx L$ the filament is termed semiflexible, coiling and knotting are not present but thermal fluctuations can induce bending Kroy and Frey (1996); if $\lambda_{\mathrm{p}}\gg L$ the filament is stiff, and it appears as a stiff rod. The persistence length can be measured from the fluctuations of the filament. Under thermal fluctuations, the tangent $\bm{t}$ to different parts of the filament performs a random walk. The correlation between two different points will then exhibit an exponential decay $\langle\bm{t}(s+\Delta s)\bm{t}(s)\rangle=e^{-\Delta s/\lambda_{\mathrm{p}}}$.
Biofilament Shear modulus $G$ Persistence length $\lambda_{\mathrm{p}}$ Young’s modulus $E$ Bending stiffness $\beta_{\mathrm{f}}$
units Pa $\mu$m kPa N m2
[…]
Keratin $6\times 10^{3}-10^{8}$ Fudge et al. (2003); McKittrick et al. (2012)
[…]
Cellulose $138\times 10^{6}$ Nishino et al. (1995)
Bacterial flagella $(0.1-380)\times 10^{6}$ Shen et al. (2022) $(0.1-100)\times 10^{-23}$ Shen et al. (2022)
Motile cilia $(0.01-120)\times 10^{6}$(†) $(0.3-6)\times 10^{-21}$ Okuno and Hiramoto (1979); Ishijima and Hiramoto (1994)
Filamentous cyanobacteria $(0.4-11.2)\times 10^{3}$ Faluweki and Goehring (2022); Boal and Ng (2010) $(16-46)\times 10^{3}$ Faluweki and Goehring (2022) $(1-20)^{-17}$ Faluweki and Goehring (2022); Kurjahn et al. (2024a)
C. elegans (adult) $(80-140)$(*)Backholm et al. (2013) $\sim 10^{-13}$ Backholm et al. (2013)
Eels and snakes $100-1000$ Guo and Mahadevan (2008); Zhang et al. (2021) $10^{-2}-10^{-3}$ Long Jr (1998)
[…]
The cytoskeleton is highly dynamic, because its constituents can be assembled and disassembled as the need arises due to permanent energy dissipation. Cytoskeletal filaments are grouped into three main classes: microtubules, actin filaments, and intermediate filaments Fletcher and Mullins (2010), see Fig. 2. The main differences among these three types of filaments are their mechanical stiffness, polarity, and the molecular motors associated with them Fletcher and Mullins (2010). We now discuss these three classes of biofilaments in turn.
$\bullet$ Microtubules are hollow and polar tubular structures, which grow by the addition of tubulin dimers ($8$ nm in length) bound to GTP (guanosine-5’-triphosphate) molecules Brouhard and Rice (2018); Zhou et al. (2023). Microtubules are the stiffest cytoskeletal filaments, see Table 1. They are formed by $\alpha$- and $\beta$-tubulin protein dimers. The protein dimers polymerize into linear chains, called protofilaments, which arranged as a hollow cylinder constitute a microtubule; most eukaryotic cells contain microtubules with 13 protofilaments, characterised by an internal diameter of about $15$ nm and an external diameter of about $25$ nm Boal (2012); Dustin (2012). Microtubules perform crucial functions in the life of a cell. During cell division, microtubule nucleate from the centrioles that organize the mitotic spindle Azimzadeh and Marshall (2010). Microtubules are also the building blocks of eukaryotic cilia and flagella Grimstone and Klug (1966); Amos and Klug (1974).
The considerable stiffness of microtubules is reflected in a persistent length $\lambda_{\mathrm{p}}$ of about 5 mm, which can span the entire length of an animal cell Fletcher and Mullins (2010), and a Young’s modulus ranging from one to 50 GPa, depending on their length, or on the formation of network structures Gittes et al. (1993); Sharma and Vershinin (2020), see Table 1. Classic experiments employed thermal fluctuations to measure the flexural rigidity of microtubules $\beta_{\mathrm{f}}=2.2\times 10^{-23}$ N m2 Gittes et al. (1993).
Microtubules are highly dynamic systems, where the cell-regulated nonequilibrium processeses of polymerisation and depolymerisation Dogterom and Leibler (1993); Fygenson et al. (1994) of a microtubule coexist (known as ‘dynamic instability’ Mitchison and Kirschner (1984)) and extend to a large fraction of the microtubules in a cell. The tau protein (short for ‘tubulin associated unit’) maintain the stability of microtubules in the axons of neurons Weingarten et al. (1975), however, when hyperphosphorylation occurs, the microtubule disintegrate and the tau proteins form aggregates (neurofibrillary tangles, which are tangles of tau filaments) that are associated with neurodegenerative diseases Palmqvist et al. (2020); Chen et al. (2024).
A single microtubule never reaches a steady state length but persists in prolonged states of polymerisation and depolymerisation that interconvert infrequently Desai and Mitchison (1997). The cell uses the free energy of GTP hydrolysis to drive the nonequilibrium polymerisation; this free energy is $\approx 7.5$ kcal mol-1 under standard conditions, or $\approx 12.5$ kcal mol-1 in vivo Lehninger et al. (1993); Desai and Mitchison (1997).
Recent measurements have shown that as microtubules grow or shrink, they undergo conformational changes. The GTP-tubulin dimers are initially curved at an angle of $\approx 12^{\circ}$; after binding they straighten but exert a stress on the rest of the filament structure; finally, after GTP hydrolysis the $\alpha$-tubulin shortens by $3$ and releases the stress, which can be harnessed to perform work Coue et al. (1991); Koshland et al. (1988); Brouhard and Rice (2018). These conformational changes in the dimers generate long-range interactions along the microtubules that modulate their growth Brouhard and Rice (2018).
Perhaps the most important example of microtubule self-organisation is the mitotic spindle, which partitions the set of chromosomes from a cell into two copies for the daughter cells Wittmann et al. (2001); Gadde and Heald (2004). The polarity of microtubules is essential in this process, as each half of the spindle contains uniformly oriented microtubules, with their minus-ends at the pole (towards the centrosome) and their plus-ends extending away. The initial growth of microtubules around centrosomes is based on their dynamic instability, with random phases of growth and shrinkage probing the cytoplasm for kinetochores (the so-called ‘search-and-capture’ model Kirschner and Mitchison (1986)). That many classes of kinesins play a role in the spindle formation is evidence that the mitotic spindle is a nonequilibrium self-organised structure Gadde and Heald (2004).
$\bullet$ Actin filaments are composed of (globular) G-actin monomers which readily assemble into long strings called F-actin (“F” for filamentous) of about 8 nm in width Boal (2012). Actin filaments are about 300 times less stiff than microtubules, with $\beta_{\mathrm{f}}=7.3\times 10^{-26}$ N m2 Gittes et al. (1993), a persistence length $\lambda_{\mathrm{p}}\approx 18~{}\mu$m Gittes et al. (1993), and with a Young’s modulus $E\simeq(1-3)$ GPa, which is comparable with Young’s modulus for bone ($9$ GPa) Mofrad (2009). Interestingly, actin is less resistant to bending than microtubules only because it is thinner, and not because it is more compliant than tubulin Gittes et al. (1993). It is also worth comparing the Young’s moduli of actin and tubulin. The measurements of $\beta_{\mathrm{f}}$ for a microtubule implies a Young’s modulus $E\sim 1.2$ GPa, while for actin $E\sim 2.6$ GPa Gittes et al. (1993). This is comparable with $E\sim 2-8$ GPa of silk (Bombyx mori) Lepore et al. (2016); Greco et al. (2022); Wainwright (1982), or with $E\sim 2.5$ GPa of keratin (from birds’ feathers) Bonser and Purslow (1995). These are all proteins employed by the cells to generate tensile strength Gittes et al. (1993), so it is not so surprising after all that filaments with similar Young’s modulus evolved for the cytoskeleton.
Both actin filaments and microtubules are polar filaments, that is, they have a positive and a negative end due to an asymmetry at the molecular level. Hence, they can function as train tracks along which molecular motors move. Dynein and kinesin are the motor proteins associated to microtubules; by means of ATP hydrolysis they can transport vesicles, mRNA, and other proteins. Myosins are the motor proteins associated to actin. They are a superfamily Hartman and Spudich (2012): at least 35 classes of motor proteins associated to actin filaments are known Odronitz and Kollmar (2007) and used for vesicle transport; they are responsible for muscle contraction in myocytes Sweeney and Houdusse (2010), and are macromolecular filaments. Some members of this superfamily, the non-muscle myosin II (NM II) are even involved in cell adhesion and migration Vicente-Manzanares et al. (2009).
Figure 2: (a) Endothelial cells from cow lung with two components of the cytoskeleton: microtubules (green) and actin filaments (red) Tina Carvalho (2011). (b) Confocal microscope image of a mouse 3T3 fibroblast cell. The nucleus has been stained blue, and two components of the cytoskeleton, actin microfilaments and intermediate filaments, are stained red and green respectively David Becker (2011).
Actin is present in high concentrations in the actomyosin cortex of animal cells, a layer attached to the inner surface of the cell membrane, and its function is to generate tension and regulate force and cell shape Étienne et al. (2015); Saha et al. (2016); Kumar (2021). Actin is the main constituent of lamellipodia, two-dimensional structures responsible for movement of e.g. epithelial cells Cardamone et al. (2011), and is involved in cell adhesion Schwarz and Safran (2013); Garcia et al. (2015). The versatility of actin is but an example of the fact that the molecular organisation of biofilaments is extremely flexible; as a further example, amyloid fibrils can change bending rigidity by four orders of magnitude due to varying backbone interactions in the cross-$\beta$ structure Knowles et al. (2007).
Actin plays a major role during embryonic development, as the cells have to undergo a specific sequence of physical and biological changes Heisenberg and Bellaïche (2013). Physically, cell cortical forces, cell adhesion and motility together produce self-organisation both at the tissue level and within single cells, e.g., reorientation of the polarity plane Aigouy et al. (2010). Another physical factor involved in tissue morphogenesis is actin-myoosin flows (or cortical flow), that is, a cell-scale internal flow of actin components of the cortex Bray and White (1988). Various aspects are at play here: the character of the actin-myosin flow dynamics (pulsatile versus continuous) and direction (centripetal or anisotropic) influence force generation; additionally, to transmit forces to other cells, this flow needs to be coupled to adhesion complexes at the cell membrane Heisenberg and Bellaïche (2013). Although much has been understood about force generation in cells and tissue, there are still many open questions about how the exquisitely complex self-organised process is controlled by the cell.
$\bullet$ Intermediate filaments (IFs) are the most flexible of the three classes of cytoskeletal filaments, but they are also a very diverse class of filaments Köster et al. (2015). Estimates of their persistence length range from $\lambda_{p}=415\pm 45$ nm for human wild-type vimentin Schopferer et al. (2009) to $\lambda_{p}=2.1\pm 0.1~{}\mu$m Nöding and Köster (2012). There are six classes of IFs Parry and Steinert (1992), which include, e.g., keratin, desmin, vimentin, and nuclear lamin. Structurally, IFs exhibit a hierarchical structure Chernyatina et al. (2015). The alpha-helical domains form a coiled-coil structural motif, composed of two parallel in-register chains. Two chains aggregate in parallel to generate an IF molecule, and then two such molecules assemble in an antiparallel fashion to form a four-chain structural unit Parry and Steinert (1992). IF proteins are highly charged, thus representing versatile polyampholytes with multiple functions due to their strong self-assembling ability Herrmann and Aebi (2016). IF is built as a complex network of ionic interactions formed by the acidic rod and the basic head domain, and, unlike microtubules and actin filaments, IFs are non-polar, and lack enzymatic activity Herrmann and Aebi (2016). IFs form networks of filaments with important mechanical properties: upon being mechanically stressed, the network becomes stiffer Janmey et al. (1991); Wen and Janmey (2011); Herrmann and Aebi (2016).
Cytoskeletal networks and viscoelastic properties. The viscoelastic properties of cytoskeletal filaments are closely linked to their biologic function Mofrad (2009), because their extreme aspect ratio and rod-like structure dominate the rheological behaviour of the cell, and changes in their structure may cause gel-sol transitions observed when cells are activated or begin to move Janmey et al. (1991). Measurements of the storage modulus $G^{\prime}$ for F-actin and fibrin suggest that the molecular basis of the viscoelasticity of these two types of networks is similar and may be approximated by the steric hindrance of diffusion of long, interpenetrated filamentous polymers Janmey et al. (1991). While the viscoelastic measurements of these biofilaments can be explained with isotropic models, there is ample evidence that microtubules and F-actin undergo a transition from isotropic to liquid crystalline solutions under the influence of shear stresses Kerst et al. (1990), other macromolecules Suzuki et al. (1989), or even spontaneously Hitt et al. (1990); Viamontes et al. (2006); Hess and Ross (2017), because of the strong thermodynamic drive for long rod-like filaments to form liquid-crystalline arrays Onsager (1949); Flory (1956).
Cytoskeletal filaments play a fundamental role in the mechanical properties and the biological function of the cell. This functional importance is due to not just the physical properties of isolated filaments, but, perhaps more crucially, to their structural organisation. Most cytoskeletal filaments form networks, and by controlling their architecture and connectivity cells achieve remarkable material properties Wagner et al. (2006); Gardel et al. (2008).
Actin filaments are assembled by actin-binding proteins into branched, cross-linked networks; the list of such proteins is vast but includes, e.g., $\alpha$-actinin, $\beta$-spectrin, filamin, and fimbrin. These complex, dynamically-reassembling, dendritic networks generate force at the cell periphery that creates membrane protrusions (lamellipodia), that in turn produce the cell’s directed motility, essential in processes such as embryonic development, wound healing, and the immune response Pollard and Borisy (2003).
The networks assembled with semi-flexible cytoskeletal filaments have unusual viscoelastic properties Pegoraro et al. (2017); Storm et al. (2005); Wagner et al. (2006); Gardel et al. (2008); Chaubet et al. (2020); Lorenz and Köster (2022). Upon increasing stress, networks of actin filaments show stress stiffening behaviour up to a critical stress of $\sigma_{c}\sim 270$ Pa, after which the networks enters a stress-softening regime that is reversible, that is, if the stress is reduced no hysteresis is observed Chaudhuri et al. (2007). A microscopic mechanism that explains such behaviour is based on the initial resistance to stretching due to entropic effects in the biofilaments, leading to the stress-stiffening regime. Above $\sigma_{c}$, some filaments under compressive stresses will buckle, which decreases the elasticity of the network and leads to the stress-softening regime Chaudhuri et al. (2007). This is at variance with the behaviour of flexible polymers networks, which exhibit stress softening only due to failure and is not reversible Mofrad (2009). Furthermore, cytoskeletal filaments show negative normal stresses, that is, when sheared they tend to pull inward Janmey et al. (2007). On the other hand, flexible-polymer networks exhibit positive normal stress. These nonlinear viscoelastic features of cytoskeletal filaments contribute to generate the structural integrity and strength of the cell.
Biofilaments exist also outside of the cell. In animal cells, collagen is the most important filament found in the extracellular matrix Sun (2021). There are many types of collagen molecules, but e.g. in humans more than $90\%$ are type I collagen, found in skin and tendons. Type I collagen forms fibers in a hierarchical structures. The amino acids chains form together a triple helix, called tropocollagen, with a diameter of $\approx 1.5$ nm. Multiple tropocollagens bundle together into protofibrils (diameter $\approx 6-25$ nm). Multiple protofibrils further bundle together and form collagen fibril (up to $1\mu$m in length, $\approx 100$ nm in diameter). The fibrils can then form thicker fibers, which self-organize into a disordered network, the collagen matrix Sun (2021). The collagen matrix exhibits nonlinear elasticity, with an elastic modulus depending on concentration, pH, and ionic strength, and ranging from few Pa up to hundreds of MPa Sheu et al. (2001); Dutov et al. (2016).
In plant cells, the most abundant and important filament is cellulose. This is a linear polysaccharide chain of glucose monomers. The numerous hydrogen bonds within and between chains gives them considerable stiffness and generates robust fibrils, which are used in the cell walls of plant cells. Depending on its microstructure, the Young’s modulus of cellulose ranges from $\approx 25$ GPa Eichhorn and Young (2001) (for microcrystalline cellulose) to $114$ GPa (for bacterial cellulose) Hsieh et al. (2008) or the classic measurement $138$ GPa Nishino et al. (1995).
A review of cell biopolymers cannot be complete without mentioning nucleic acids: DNA and RNA. The physics of the DNA helix, or generally of nucleic acids, is surprisingly complex Strzelecka et al. (1988), and is inspiring a number of applications in nanotechnology Ramsay (1998); Seeman (2010); Saccà and Niemeyer (2012); Seeman and Sleiman (2017); Praetorius et al. (2017); Engelen and Dietz (2021); but, unfortunately, we do not have the space to address it here. There are, however, numerous excellent reviews on DNA’s and RNA’s structure and biochemical interactions Kornyshev et al. (2007); Wong and Pollack (2010); Teif and Bohinc (2011); Travers and Muskhelishvili (2015); Michieletto and Gilbert (2019) and also on application of topology to DNA Chevizovich et al. (2020); Michieletto (2024)
III Cellular appendages
In a letter sent to the Royal Society of London in 1676, van Leeuwenhoek reported the discovery of protozoa and the observation of cilia and flagella. He wrote: “I also discovered a second sort of animalcules, whose figure was an oval, … provided with diverse incredibly thin little feet, or little legs [cilia], which were moved very nimbly …, and wherewith they brought off incredibly quick motions” (cf. translation of letter 18 Dobell (1958); Haimo and Rosenbaum (1981)).
Today, our language has changed a bit, and we describe cilia and flagella as slender, hair-like appendages that extend from the main cell body and perform diverse functions, such as enabling cell movement Beeby et al. (2020); Bondoc-Naumovitz et al. (2023), generating flows Eichele et al. (2020); Huang and Choma (2015); Ling et al. (2024), sensing environmental stimuli Marshall and Nonaka (2006); R. Ferreira et al. (2019) and attachment to surfaces Ryu and Matsudaira (2010); Jarrell et al. (2013). While there are a myriad of different filamentous appendages found in biology, here our focus is on those most widely studied, while also aiming to capture the diversity of their structures, properties, mechanisms and functionalities.
Across the tree of life, the most well studied cellular appendages are those used for motility, in particular the flagella in prokaryotes and cilia in eukaryotes. These filaments whip, wave, or rotate through the action of molecular motors that transform the cell’s chemical energy into mechanical energy Beeby et al. (2020); Gilpin et al. (2020). The resulting non-reciprocal motion enables net propulsion within the cell’s microscale, viscosity-dominated environment Purcell (1977); Lauga and Powers (2009).
Box III | Drag anisotropy Drag anisotropy refers to the directional dependence of drag forces acting on an object as it moves through a medium. Due to their elongated form, filaments experience less hydrodynamic drag when pulled along their length than when pulled perpendicular to it. Drag anisotropy can be quantified as the ratio $\xi=\frac{\zeta_{\perp}}{\zeta_{\parallel}},$ (3) where $\zeta_{\perp}$ and $\zeta_{\parallel}$ are the friction coefficients for motion perpendicular and parallel to the object’s length. Typically, for filaments $\zeta_{\perp}\approx 2\zeta_{\parallel}$. The total drag force density (drag per unit length) opposing a filament’s motion at low Reynolds number (i.e. viscosity-dominated regime) is given by $\textbf{F}_{\text{d}}=-(\zeta_{\parallel}\textbf{v}_{\parallel}+\zeta_{\perp}\textbf{v}_{\perp})$, where $\textbf{v}_{\parallel}$ and $\textbf{v}_{\perp}$ are the projections of the local velocity into directions parallel and perpendicular to the segment Lauga and Powers (2009); Hancock (1953).
A low Reynolds number world. At the microscale in which these filaments operate, inertia is insignificant and viscous forces dominate. In fluid dynamics, this is the world of ‘Low Reynolds number’. The Reynolds number describes the ratio of inertial to viscous forces and is defined as $\text{Re}=\frac{\rho UL}{\eta}$, where $\rho$ is the fluid density, $\eta$ its viscosity, and $U$ and $L$ are characteristic velocity and length scales, respectively. In the world of a planktonic cell, $\textrm{Re}\approx 10^{-5}-10^{4}$, thus, in this viscosity-dominated environment, cells overcome and even exploit drag to generate motion Purcell (1977); Lauga and Powers (2009). Cells can achieve this due to the anisotropic drag (defined in Box III) of slender appendages, which allows propulsive forces to be generated perpendicular to the local velocity of the filament Lauga and Powers (2009). Additionally, drag-based thrust relies on the ability of filaments to perform time-periodic deformations that are also time-irreversible motions, crucial for creating a non-zero time-averaged propulsive force Purcell (1977); Lauga and Powers (2009). This can be achieved via the rotation of helical flagella or the whip-like beating of cilia.
Multiscale dynamics. The activity of cellular filamentous appendages encompasses dynamics across multiple scales Gilpin et al. (2020), offering physicists many opportunities to develop frameworks and provide insights into how these dynamics integrate and interact across scales. Nanoscale molecular motors drive shape changes and motions of individual filaments, while these shape changes drive fluid flows Gilpin et al. (2020); Beeby et al. (2020). Resistive force theory (RFT) and slender body theory (SBT) are two hydrodynamic models of fluid-structure interactions for slender bodies within a low Reynolds number environment (see reviews Lauga and Powers (2009); Bondoc-Naumovitz et al. (2023) and references therein). Both are used to estimate the propulsive forces generated by slender appendages based on their waveforms and resulting microswimmer velocities. While RFT is a simpler approximation that neglects long-range hydrodynamic interaction, SBT is more accurate but also more computationally intensive. Mechanical coupling, both via hydrodynamic interactions and the internal cell cytoskeleton, enables coordination of multiple appendages Wan and Poon (2024); Wan (2018), and motivates a range of work that considers cellular appendages as coupled oscillators Uchida et al. (2017); Bruot and Cicuta (2016). Arrays of cilia can synchronise into metachronal waves Poon et al. (2023); Meng et al. (2021); Wan and Poon (2024); Elgeti and Gompper (2013), while the bundling and unbundling dynamics of groups of bacterial flagella are key for their run-and-tumble motility behaviour Grognot and Taute (2021); Wadhwa and Berg (2022). Large numbers of cilia can also group together to form compound ciliary structures, such as the leg-like ‘cirri’ the ciliate Euplotes uses for walking Laeverenz-Schlogelhofer and Wan (2024) and the comb plates of ctenophores that comprise tens of thousands of cilia Jokura et al. (2022). The individual and collective activity of filamentous appendages drive fluid transport, self-propulsion and organism-environment interactions. Groups of microswimmers can display collective behaviours such as bacterial swarming Be’er and Ariel (2019); Wadhwa and Berg (2022) and formation of bioconvection patterns that arise from hydrodynamic instabilities Bees (2020); Schwarzendahl and Mazza (2018); Vachier and Mazza (2019); Schwarzendahl and Mazza (2019, 2022).
Figure 3: Four principles of activation in filamentous cellular appendages. (a) Rotationally actuated filaments are typically rigid helical filaments rotated at their base, such as bacterial flagella seen here in an image reproduced from Turner et al. (2000) of a fluorescently labelled Escherichia coli cell. (b) Linearly actuated bending is the mode of activation found in the cilia of eukaryotes, with typical waveforms consisting of power and recovery strokes. The microalga Chlamydomonas uses two cilia for motility, here shown as three overlaid frames from a movie of its swimming behaviour. (c) Assembly-disassembly activity involves the dynamic extension and retraction of filaments, such as in pili used to generate twitching motility in bacteria and archaea. An example cryo-electron micrograph of a negatively stained wild-type Thermus thermophilus cell shows long flexible pili filaments Neuhaus et al. (2020). (d) Pre-stressed springs respond to changes in intracellular calcium resulting in rapid movements, as seen in the time series of the contractile motion of the stalk of Vorticella Ryu and Matsudaira (2010) and the haptoneme of a haptophyte alga Nomura et al. (2019).
Considering the variety of mechanisms cells employ to actuate their filamentous appendages, these can be broadly classified into four different principles of activity: rotational actuation of passive helical propellors, linear motor actuation of flexible filaments to generate bending, assembly-disassembly dynamics, and pre-stressed springs; see Fig. 3 for the corresponding schematics and examples.
Rotationally actuated filaments. The rotation of passive filaments by cells involves a rotary complex embedded in the cell membrane at the base of the filament where stator and rotary components interact to generate torque that is transmitted to an inert rigid helical filament. There are two main examples of such filaments: the bacterial flagella and the archaea’s archaella Beeby et al. (2020). Their molecular structures differ. The bacterial flagellum is a hollow cylinder self-assembled from the flagellin protein into a a chiral structure Silverman and Simon (1977); Coombs et al. (2002). The archaea’s archaella has a complex structure Cvirkaite-Krupovic et al. (2023). However, they are both $\sim$10-20 nm wide and up to $\sim$10 $\mu$m in length. Archaella use an ATPase to generate torque via ATP hydrolysis Streif et al. (2008); Albers and Jarrell (2015), while the flagella’s rotary complex takes a more structurally intricate form and is powered by the proton motive force Berg (2003); Manson et al. (1977); Wadhwa and Berg (2022). In both cases, the torque generated by the embedded molecular machinery results in a rotation of the connected passive rigid filament that in turn acts as a propellor for cell motility Beeby et al. (2020).
Linearly actuated filaments. In contrast, cilia (sometimes referred to as eukaryotic flagella) are flexible filaments, typically $\sim$250 nm wide and up to $\sim$100 $\mu$m long, and are actuated by linear stepper motors distributed along their entire length Beeby et al. (2020). Cilia have a highly conserved structure across eukaryotes known as the axoneme, which comprises 9 microtubule doublets surrounding a central pair of microtubules Nicastro et al. (2006). Ciliary bending and twisting is powered by dynein motors that walk along the 9 outer doublets. Since the microtubules are usually fixed to the cell body at the base of the cilium, this linear forcing generates differential sliding between neighbouring doublets resulting in bending of the axoneme Lindemann and Lesich (2010); Satir et al. (2014). The mechanism by which dynein motor activity generates ciliary beating is still under debate, with several biophysical models proposed that largely fall under two categories: (1) mechanical feedback within the axoneme regulates dynein activity Lindemann and Lesich (2010); Riedel-Kruse et al. (2007); Cass and Bloomfield-Gadêlha (2023), (2) continuous dynein activity along the flexible filament results in dynamic instabilities Woodhams et al. (2022). Dynein motors are typically distributed along the whole length of the cilium, giving a high degree of flexibility in the types of waveforms that are possible to produce by coordinating the dynein activity in different ways. Compared with the rotation of a rigid filament, which only has a small set of parameters that can be controlled (i.e. speed and direction of rotation), cilia have the potential to display a much more complex and nuanced diversity of waveforms, with calcium playing a central role in the regulation of ciliary activity Eckert (1972). The ability of cilia to change their waveform is critical for enabling ciliated organisms to switch locomotor gaits and perform complex motion Wan and Goldstein (2018); Laeverenz-Schlogelhofer and Wan (2024). Although cilia are highly conserved, there are several examples of variations or additions to the standard 9+2 axoneme structure that exemplify how cilia can be adapted to perform specific functions Ginger et al. (2008). For example, the cilium of the parasite Trypanosoma brucei has a paraflagellar rod, a lattice-like filament that runs along the length of the axoneme, thought to affect axonemal beating by imposing mechanical constraints and contributing to signaling pathways Langousis and Hill (2014). Excavates, species of phagotrophic flagellates, are characterised by possessing vaned cilia within a ventral groove whose confined beating is thought to be particularly efficient for generating feeding currents Suzuki-Tellier et al. (2024).
Assembly-disassembly dynamics. An entirely different principle of activity is that seen in the type IV pili of bacteria and archaea, in which cycles of extension, surface binding and retraction enable cells to perform functions such as twitching motility, surface sensing and DNA uptake Craig et al. (2019). This activity is enabled by ATP hydrolysis, which powers the rapid assembly and disassembly of pilin subunits Ellison et al. (2022).
Pre-stressed springs. Lastly, pre-stressed springs composed of filaments and tubules power many of the fastest cellular movements Mahadevan and Matsudaira (2000). There are two main types: a microtubule-based system found in some microalgae (i.e. the haptoneme in haptophytes) Nomura et al. (2019) and a myoneme based system found in some ciliates (e.g. the spasmoneme of vorticella) Ryu et al. (2017); Floyd et al. (2023). In both cases, rapid coiling is thought to be driven by an increase in Ca2+ concentration, which changes the energetically preferred conformation of the filament from elongated to coiled Kawachi and Inouye (1994); Floyd et al. (2023). For the haptoneme, some evidence suggests that this is driven by conformational changes in microtubule associated proteins that bind with Ca2+ Nomura et al. (2019), while for the spasmoneme it is hypothesised to be a combined electrostatic and entropic effect Mahadevan and Matsudaira (2000).
Elastodynamic properties. A key mechanical parameter for characterising the movements of cellular appendages is the bending stiffness (see Box II), which can be estimated from the deformation induced by forces and torques applied using techniques such as optical tweezers Darnton and Berg (2007), microneedles Xu et al. (2016), magnetic beads Hill et al. (2010) or microfluidic flow devices Shen et al. (2022). Using such approaches the bending stiffness for the bacterial flagella has been estimated to be $1-1000$ pN $\mu$m2 (see table S1 in Shen et al. (2022)), while estimates for eukaryotic motile cilia fall within the 300-6000 pN $\mu$m2 range Okuno and Hiramoto (1979); Hill et al. (2010); Xu et al. (2016); Ishijima and Hiramoto (1994). The Young’s modulus $E$ is generally estimated indirectly from experimentally measured values of the bending stiffness. This requires approximating the second area moment $I$, which is challenging when a filament is not easily defined as a solid or hollow cylinder, such as in the case for cilia. For motile cilia, two values are commonly used: $I=1.3\times 10^{-31}$ m4 Holwill (1965) and $I=3\times 10^{-29}$ m4 Baba (1972), the latter assumed more relevant for tighter connections between the microtubulues of the axoneme Okuno and Hiramoto (1979). This gives a large range of estimates $E\sim(0.01-120)$ GPa.
Appendages as cellular antennae. In addition to their role in motility and fluid transport, appendages can also play important roles in sensory reception and signal transduction that benefit from the geometric and mechanical properties of their filamentous form Marshall and Nonaka (2006); Hickey et al. (2021); R. Ferreira et al. (2019); Craig et al. (2019). Cilia often possess many sensory structures such as chemical receptors and mechanically-sensitive ion channels that enable them to act as the cell’s antenna Marshall and Nonaka (2006); Bloodgood (2010). Filamentous appendages can extend far beyond the cell surface, reaching more well-mixed and therefore more ‘representative’ regions of the surrounding environment, making them effective as chemosensory probes Marshall and Nonaka (2006). Additionally, their high aspect-ratio and high surface area to volume ratio, improves their capture rate and chemosensory effectiveness Hickey et al. (2021). Their elongated shape also means that filamentous appendages can be considered as extended cantilevers, facilitating their sensitivity to mechanical stimuli R. Ferreira et al. (2019).
IV Filamentous microbes
Figure 4: Examples of cyanobacteria self-organisation. (a) Many cyanobacteria species grow into filaments made out of a chain of identical cells Culture Collection of Algae, SAG . (b) These filaments are capable of organizing into intricate structures like reticulate mats Faluweki and Goehring (2022), spirals Gong and Prakash (2023) (c), and even ropes Garcia-Pichel and Wojciechowski (2009) (d).
Various microbial organisms have evolved a filamentous body shape. Examples of such organisms that have attracted the interest of physicists include various filamentous cyanobacteria Faluweki et al. (2023); Faluweki and Goehring (2022); Kurjahn et al. (2024b, a); Cammann et al. (2024); Ehlers et al. (1996), filamentous phages Dogic and Fraden (1997, 2006); Purdy and Fraden (2004), and nematodes, Bilbao et al. (2013); Gart et al. (2011); Backholm et al. (2013); Niebur and Erdös (1991) that straddle the boundary between the microscopic and macroscopic world. As there are too many filamentous microorganisms to list, this section will focus on discussing more generally the implications of a filamentous body shape at the microscale for a few important classes of microorganisms.
Having a body with a large aspect-ratio like a filament comes with some advantages over microbes shaped as spheres or short rods. An elongated shape increases the chance of meeting one another in disordered environments Arguedas-Leiva et al. (2022). Thanks to a filamentous shape, microbial organisms can self-propel without appendages. Rather than using flagella or cilia to generate motility, they can bend their whole body to generate thrust thanks to the different drag coefficients that a slender body experiences in the direction tangential and perpendicular to its motion in a viscous fluid Lauga and Powers (2009); Hancock (1953) (see Box III). A filamentous shape has been observed to be a viable strategy to avoid predation by protists by simply being too long to be ingested easily Pernthaler (2005) (this approach is however futile at a certain size difference between devourer and the devoured, as is evident in the popularity of spaghetti as a human food source). The increased surface-to-volume ratio of a filamentous shape allows for easier uptake of nutrients Young (2006), but also allows stronger attachment to substrates Weiss et al. (1995) in bacteria, because of a larger contact patch, compared to a spherical one.
A filament’s most crucial mechanical property is its bending stiffness. It is known one can draw conclusions about the forces at play from purely geometrical observations (see Box II). Techniques to measure the bending stiffness $\beta_{\mathrm{f}}$ and Young’s modulus $E$ of gliding filaments with the help of microfluidics Faluweki and Goehring (2022) and mircopipette techniques Kurjahn et al. (2024a) have been developed (see Table 1).
Filamentous cyanobacteria often exist in a surface-bound but motile state, called gliding Hoiczyk (2000); Wilde and Mullineaux (2015); Tchoufag et al. (2019). The underlying mechanism that allows the generation of propulsive forces in this state is still a matter of debate McBride (2001); Read et al. (2007); Dhahri et al. (2013); Koiller et al. (2010); Halfen and Castenholz (1970), as no cyanobacteria species are known to possess flagella Wilde and Mullineaux (2015). The main mechanisms that are debated are: type IV pili extension and retraction Khayatan et al. (2015); Wilde and Mullineaux (2015), fibrils Halfen and Castenholz (1970); Read et al. (2007), ejection of a jet of extracellular slime from junctional pores Hoiczyk and Baumeister (1995, 1998); Wolgemuth and Oster (2004); Dhahri et al. (2013) (at least for some species), and acoustic streaming Ehlers et al. (1996); Koiller et al. (2010); Ehlers and Koiller (2011). By observing the self-buckling of a filament, the forces at play in its propulsion can be inferred via the bending stiffness Kurjahn et al. (2024a). Such observations combined with the independence of propulsion speed on filament length implies, that all cells within a cyanobacteria filament contribute roughly equal to its propulsion.
Cyanobacteria locomotion exhibits another interesting feature: path-tracking motion, where the body follows the path laid by the forward end of the filament Faluweki et al. (2023). This type of locomotion in worms is also known as metameric locomotion Du et al. (2022). Physically, information about a filament’s orientation persists over the time it takes a filament to traverse its own length Du et al. (2022). Although their motility is induced by polar forces, filaments themselves are seemingly nematic, with no head or tail end. In fact, cyanobacteria have been observed to reverse their direction of motion frequently, as a strategy to navigate towards more favorable conditions, and even to form communities. Wilde and Mullineaux (2015); Rosko et al. (2024); Abbaspour et al. (2023); Pfreundt et al. (2023). Given the delocalised nature of the filaments propulsive mechanism, a high degree of intercellular coordination is required to reverse the direction of motion, with the mechanisms of communication between cells not yet understood. Confined active filaments with low bending rigidity have been shown to curl up into tightly wound spirals (see Fig. 4(c)), effectively trapping themselves, greatly impeding their mobility Duman et al. (2018); Prathyusha et al. (2018); Isele-Holder et al. (2015); Lin et al. (2014). Utilising reversals is a strategy to escape this trapped state Gong and Prakash (2023).
Having a high aspect-ratio allows microbes to effectively sample changing conditions within space. While organisms with short bodies commonly have to move around to find favourable conditions, essentially sampling the environment at random to inform their movement through chemo- or phototaxis, elongated filaments can sample gradients along their body and react before they have wholly immersed themselves in an unfavourable environment. Cyanobacteria, for instance, have been observed to react to light shining only onto parts of their body Nultsch and Wenderoth (1983); Wilde and Mullineaux (2017), showing scotophobic responses, well before the whole filament loses access to light by reversing its gliding direction Kurjahn et al. (2024b).
Having an elongated body may even allow organisms to be in two places with different conditions at once. Cable bacteria are filamentous, multicellular bacteria belonging to the family Desulfobulbaceae. They can conduct electrons over centimetre scales Pfeffer et al. (2012); Marzocchi et al. (2014). Through their filamentous shape they can couple sulfide oxidation in deeper, anoxic marine sediment with oxygen reduction in surface layers Nielsen et al. (2010); Bjerg et al. (2023); Risgaard-Petersen et al. (2015). Electrical activity is also involved in a microbial feat of engineering: some electroactive bacteria produce nanowires to transfer electrons to extracellular electron acceptors Reguera et al. (2005). These are, e.g.,the Geobacter sulfurreducens or the Shewanella oneidensis MR-1 bacteria El-Naggar et al. (2010); Pirbadian et al. (2014); Schkolnik et al. (2015); Logan et al. (2019).
Filamentous microbes can explore spaces inaccessible to plumper organisms with the same volume. Semiflexible active filaments have been shown to effectively explore porous environments Mokhtari and Zippelius (2019). Many soil dwelling microbes are filamentous in shape, in addition to their ability to explore a porous environment it has been speculated that their shape may enable them to fix their position in space, by wrapping around soil particles Young (2006), allowing them to resist being washed away by percolating water. Entangled filamentous microbes have been observed to stabilize their habitats and protect them against erosion Kurtz Jr and Netoff (2001); Pluis (1994). Furthermore, nonmotile cells exploring a polymeric fluid can form long cables as they proliferate Corte et al. (2025).
Their filamentous nature allows filamentous organisms such as cyanobacteria to organise into structures much larger than an individual filament Faluweki et al. (2023); Cammann et al. (2024); Shepard and Sumner (2010); Tamulonis and Kaandorp (2014). Their filamentous shape allows them to glide over one another without losing contact with the substrate. The dynamics of gliding filaments result in the formation of a network-like reticulate pattern of filament bundles Faluweki et al. (2023); Cammann et al. (2024); Tamulonis and Kaandorp (2014) (see Fig. 4(b)). Under different conditions, filaments have been observed to form aster-like aggregates Pfreundt et al. (2023). These relatively simple, mostly two-dimensional structures may serve as a starting point for filaments to organize themselves into even more intricate three-dimensional structures, ranging from complex patterned biomats Shepard and Sumner (2010); Cuadrado and Pan (2018) to the filaments weaving themselves into ropes Garcia-Pichel and Wojciechowski (2009) (see Fig. 4(d)), bridging large distances. The physical mechanisms behind this level of self-organisation and how the mechanical properties of these structures compare to that of a single filaments is yet to be understood, but can shed light on biofilm formation Mazza (2016); Sauer et al. (2022). It is plausible to presume that such complex assemblys could exhibit unusual rheological properties, just like those observed in cytosceletal networks (see Sec. II).
V Undulatory locomotion at the macroscopic scale
At the organismal scale, animals with a slender and deformable body shape are also ‘active filaments’, and in some cases the analysis and framework presented in the previous sections could still apply. Diverse animals locomote by propagating undulatory waves. These body deformations are used to propel the organisms forward, although the actuation mechanisms may be very different across species and spatial scales (Figure 5). Well-studied examples include crawling worms Costa et al. (2024); Stephens et al. (2008), swimming eels Gillis (1998), and gliding snakes Jayne (2020). The study of both limbed and limbless movement has paved the way for understanding the origins and principles of neuro-mechanical control in diverse organisms, particularly the interactions between neural patterning and muscular activity Grillner et al. (1991). Undulatory or sinusoidal locomotion strategies evolved across different phyla, often according to a basic neural circuit design involving mutually inhibitory halves with the left and right sides deforming alternately Katz (2016). These insights into the neural basis of behaviour have led to the development of novel bio-inspired mechanisms for actuating synthetic systems including robots Ramdya and Ijspeert (2023).
Since wave-like deformations are inherently non-reciprocal, they form the basis of diverse motility strategies through different media. In some cases, despite the large difference in scale compared to microscopic organisms such as flagellates and cyanobacteria (see Sec. III and IV), the undulatory locomotion of macroscopic animals such as nematodes and snakes can still be considered to operate in an overdamped regime where inertia is negligible, and where friction (instead of fluid viscosity) dominates. This is particularly true of locomotion in sand or granular media, where a resistive force theory based on drag anisotropy (see Box III) has been derived empirically, inspired by the movement of slender bodies through viscous fluids Hatton and Choset (2011); Zhang and Goldman (2014).
Figure 5: Undulatory locomotion of active filaments across scales. (a) For C. elegans crawling movement on agar Padmanabhan et al. (2012). (b) Sidewinding locomotion of a snake, and (c) the timetrace of tracked points along the body of the snake as it moves on sand Rieser et al. (2021). (d) Vinegar eels self-organise into metachronal waves Peshkov et al. (2022). (e) Topological entanglement in worm blobs Patil et al. (2023). (f) Bio-inspired snake-like robots with CPG-like control modules Wu and Ma (2010).
At the individual organism scale, distinct undulatory behaviours can arise from the same underlying structure and neural circuitry. A given body shape can be decomposed into a small number of principal components or oscillation modes, reflecting the intrinsically low-dimensional nature of the posture space, and allowing diverse undulatory shape modes to be reconstructed with only a small number of parameters Yemini et al. (2013). In the example of C. elegans crawling, its motion in 2D is stereotyped and well-captured by just six basis modes, or eigenworms Stephens et al. (2008); Costa et al. (2024). Frequency and amplitude thus uniquely define a gait, which in turn determines the distance traversed by the organism in one cycle. Notions such as geometric phase can help understand and predict the mapping between these gaits and eventual displacement in the lab frame Rieser et al. (2024). These dynamics are not totally unconstrained however, undulatory locomotion across swimming sperm, nematodes, and even fish in overdamped environments may be subject to a universal dispersion relationship Pierce et al. (2024); Sánchez-Rodríguez et al. (2023). Thus, active filaments can achieve highly reconfigurable dynamics with minimal neuromechanical coupling.
Collectively, filamentous organisms also engage in novel behaviours that may not be accessible by individuals. At sufficiently high densities, these filaments self-assemble to form adaptive active materials with both solid-like and liquid-like properties, that are capable of emergent functions such as directed locomotion or smart navigation. Macroscopic ‘blobs’ of the 1 cm-long California blackworm Lumbriculus variegatus undergo dynamic collective movements to avoid desiccation or thermal stress without the need for any centralised control Ozkan-Aydin et al. (2021). The entanglement, or ‘braiding’, process enables fast and reversible tangling and untangling of worm collectives. Intriguingly, L. variegatus blobs can take minutes to assemble but only milliseconds to disentangle. A recent study proposed an explanation for this phenomenon by experimentally imaging the 3D trajectories of the worms and mapping the dynamics onto a 3D filament model of Kirchhoff rods Patil et al. (2023). The unique, activity-dependent rheology of entangled blobs of Tubifex tubifex worms was also studied and shown to be shear-thinning, but with a different scaling compared to polymer melts Deblais et al. (2020). These properties can even be harnessed to enable activity-dependent separation (chromatography) of worm mixtures Heeremans et al. (2022). In another free-swimming nematode Turbatrix aceti (vingar eels), concentrated suspensions of these organisms display large-scale metachronal waves surprisingly similar to those found in ensembles of cilia and flagella Peshkov et al. (2022); Quillen et al. (2021). Understanding how these deformable filamentous bodies actively control the topology and rheology of the collective could inspire novel designs of functionally tunable synthetic materials Deblais et al. (2023).
More generally, undulatory or anguilliform locomotion in elongated bodies allows for direct interactions between an organism and the surrounding medium, which can yield novel principles of biomimetic/bioinspired locomotion reliant on embodied design with few mechanical components. Self-propulsion can also be easily adapted for movement through different environments, for example for snakes the same basic locomotion mechanism can be equally effective for locomotion on land as well as in water Taylor (1952); Stin et al. (2024). In more extreme cases, ‘flying’ snakes from the genus Chrysopelea use body undulations to generate lift to glide stably over several metres through the air Yeaton et al. (2020).
From a control theoretic perspective, undulatory locomotion is universal perhaps because it reduces the need for more complex neural control mechanisms and can be achieved simply by coordinating alternating waves of (usually muscle) contraction and relaxation. In some cases, oscillations emerge spontaneously as a result of instabilities, for example dynamics reminiscent of ciliary beating have been observed in macroscopic chains of self-propelled microrobots, with the transition mapped to a supercritical Hopf bifurcation Zheng et al. (2023). Indeed substrate-attached, microscopic elastic filaments forced at one end also exhibit oscillations in different regimes De Canio et al. (2017); Bayly and Dutcher (2016); Man et al. (2020). The propagation of waves thus provides one of the simplest possible appendage or body actuation mechanisms, and naturally breaks time-reversal symmetry (particularly relevant for inertia-free swimming). This results in highly adaptable and flexible control of movement where both the speed and direction of travel can be readily changed in response to stimuli.
VI Roadmap & Key Open Questions
Filamentous structures are ubiquitous in biology from cytoskeletal filaments to propulsive appendages to whole organisms. Despite spanning several orders of magnitude in scale and complexity, surprising parallels can be observed in how biological form translates into physiological function in these different contexts. In this review, we have sought to outline the core physical principles that govern these relationships and common frameworks for studying filamentous systems. As cutting-edge experimental technologies continue to deliver ever finer resolution of the molecular composition and architecture of the constituent parts of organelles, cells, and organisms, there is increasing need to integrate these insights with functional assays and biophysical models. At the same time, we should also be mindful that in many cases the natural context or habitat in which organisms live could also contribute to any observed phenomena Franks et al. (2016). Therefore, there is also a growing need to study and understand the behaviour of living systems ‘in the wild’, not just under controlled laboratory settings.
Critical to many modelling and simulation approaches is detailed knowledge of the emergent material properties of the many types of filaments. Yet such measurements are not always available as they often rely on access to specialised technology, such as high-resolution imaging, atomic force microscopes, micropipette force sensors Backholm et al. (2013), or can only be performed on certain filaments, or in non-living systems or systems that have been rendered passive. This apparent disconnect requires more holistic, physics-inspired understanding of the context-dependent material properties of a filament and filamentous network, particularly due to their out-of-equilibrium activity. As a concrete example, the measurements of physical parameters such as Young’s modulus $E$ or the bending stiffness $\beta_{\mathrm{f}}$ of cytoskeletal filaments as shown in Table 1 strongly depends on the chemical and physical context.
Recently, theories of active polymers Winkler and Gompper (2020) and active nematics Doostmohammadi et al. (2018) have identified interesting collective behaviour (such as bands, asters, and topological defects Liverpool et al. (2001); Duman et al. (2018); Bianco et al. (2018); Prathyusha et al. (2018); Vliegenthart et al. (2020)) that reproduce experiments with purified systems of microtubules and molecular motors Nédélec et al. (1997); Sanchez et al. (2012); Kruse et al. (2004); Sumino et al. (2012); Huber et al. (2018); Sciortino and Bausch (2021). While these systems shed light on the nonequilibrium dynamics of cytoskeletal filaments, the challenge that lies ahead is to develop these theories towards more realistic conditions like those inside the cell, which is a 3D crowded environment. Thus, these theoretical frameworks should be developed in tandem with experimental innovations. Identifying appropriate dimensionless quantities from filament properties, like the ones in Table 1, will enable the development of theories describing common phenomena in the collective dynamics of filaments. Doing so will allow the development of experimentally convenient proxy systems. Thus, efforts in characterising filaments, completing and expanding the data in Table 1 is required to develop a unified understanding of filament dynamics across scales.
The study of active or self-propelled filaments in biology have led to advancement of new mathematical and physical theories, notably in the areas of topological braiding and entanglement, nonequilibrium statistical physics, and extreme mechanics such as supercoiling. However, one constraint in understanding filamentous systems is the practical challenge of acquiring good resolution data, and the lack of efficient computational tools to track filaments, particularly from 3D image stacks. There is also the open question of which metrics are most appropriate to describe the deformation of filaments and their collective organisation. Thus, devising universally-applicable, relatively system-independent measures for such systems requires further innovation.
We speculate that, because living systems apparently increase entropy production Martyushev and Seleznev (2006); Davies et al. (2013) while producing more complex structures (Fig. 1) during their evolution, a generalised form of entropy might be a useful quantity in the search of system-independent measures. Because configurational degrees of freedom (positions, orientations) can be measured, it is possible to define structural complexity (or entropy) and use it to characterize living systems. Consider a function $\psi(\bm{x})$ representing the configuration of a system, for example, the atomic positions in a protein, the orientation of filaments in a network, or the muscle fibers in a vertebrate. The complexity of an ensemble can be quantified by means of a generalised Shannon–Boltzmann entropy Adami et al. (2000); Adami (2002); Bonchev and Buck (2005) ${H}=-\int D\psi\,P(\psi(\bm{x}))\ln P(\psi(\bm{x}))$, where $\int D\psi$ represents a path-integral sum over all possible configurations. We note that elongated shapes have five degrees of freedom (compare with the three degrees of freedom for a spherically symmetric object); furthermore, a flexible elongated shape has 5 degrees of freedom per unit length. Such explosion of degrees of freedom might constitute an entropic incentive for biological systems to evolve and use biological filaments.
It is also instructive to ask why filaments are a recurring motif across living systems. Not only do single units admit complex shape deformations, filaments can also easily self-assemble to form higher-order structures, or be reconfigured into networks of increasing complexity. These active physical structures have the capacity to perform complex functions without sophisticated control or centralised computation Wang et al. (2023); Moreau et al. (2024). Future applications, particularly in the realm of synthetic biology or bioengineering, may seek to harness the mechanical intelligence inherent in filamentous structures, to create novel adaptive materials.